You are here

Observation

Status: 
Wykonane
Observation time: 
Friday, April 13, 2012 - 20:30 to Saturday, April 14, 2012 - 19:00
Observer: 
Instruktor: 
Is scope parked?: 
Tak
Is the dome parked?: 
Tak

Goto-targeting was off but now fixed.

A lot of humidity.

 

The photometric filter wheel is placed inside the CCD.

The spectrograph was left mounted on the telescope.